OpenStudy (anonymous):

does the equation cube root x= 1-x have a solution? use intermediate value theorem

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

to use the intermediate value theorem, make your function \[f(x)=x-\sqrt[3]{1-x}\]

OpenStudy (misty1212):

if it is negative for some value of \(x\) and positive for some other value of \(x\) since it must be zero in between by the IVT

OpenStudy (anonymous):

I dont understand.

OpenStudy (anonymous):

I graphed the equation from both sides, but Im not sure the point where it intersects.

OpenStudy (anonymous):

?

OpenStudy (misty1212):

let me look

OpenStudy (misty1212):

\[f(x)=x-\sqrt[3]{1-x}\] \[f(0)=-1\\ f(9)=7\]

OpenStudy (misty1212):

you can use other numbers if you like since it is negative somewhere and positive somewhere else, it must be zero in between

OpenStudy (misty1212):

actually you could even use \(f(2)=3\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

could i ask you another question?

OpenStudy (misty1212):

sure